The median of a trapezoid is 15 dm.

, and one base is 20 dm. Find the other base.



1
Expert's answer
2021-02-03T10:15:16-0500

Solution:

To find the median use the following formula:

median=((length of one base) + (length of the other base)) / 2.

According to the task the length of one base is 20 dm and the median is 15 dm.

So, (length of the other base)=(2*median)-(length of one base)=2*15-20=30-20=10


Answer:

The length of the other base is 10 dm.